OpenStudy (anonymous):

can someone give me a quick run down of how to factorise say x^2+x-6=0

OpenStudy (anonymous):

take the product of the last term and the coeff of the first this case -6 x 1 = -6 now break the middle term into factors of -6 x = 3x - 2x

OpenStudy (saifoo.khan):

(x - 2) ( x + 3) = 0

OpenStudy (anonymous):

goodness paul... ok we're going to generalize for almost every factorable expression this will work, but this is generally the long way: Ax^2+Bx+C=0 find two numbers where A+C=B and B=AC these 2 numbers will be called y and z. Ax^2 + yx + zx + C=0 then use grouping and follow through. x^2 + x - 6 = 0 y and z would be -2 and 3. x^2 - 2x + 3x -6 = 0 x(x-2) + 3(x-2)=0 The two quantites will always be the same. If not you did something wrong. Now you can use the distributive property. (x+3)(x-2)=0 Now you know one of the two quantities has to be equal to 0, so create 2 equations and solve for x to get your 2 possible answers.

OpenStudy (anonymous):

That took me a long time to write so I hope that makes sense...

OpenStudy (anonymous):

h m can you give me another one that i can just practice on then? im pretty sure that made sense, but i want to be certain.

OpenStudy (anonymous):

This course you're taking is weird Paul

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

Well, it just is...I guess it's more of a review course

OpenStudy (anonymous):

thats exactly what it is.

OpenStudy (anonymous):

Hey paul I can show you how to factorize step by step

OpenStudy (anonymous):

please do

OpenStudy (anonymous):

Okay, hold on.

OpenStudy (anonymous):

http://www.wiziq.com/online-class/575722-paul-class

OpenStudy (anonymous):

Paul, have you clicked the link yet?

OpenStudy (anonymous):

yeah i started the class :(

OpenStudy (anonymous):

i can hear you